OpenStudy (anonymous):

Please Help Establish!!! cos2(2u)−sin2(2u)=cos(4u)

OpenStudy (anonymous):

cos^2(2u)−sin^2(2u)=cos(4u)

OpenStudy (anonymous):

\[\cos^2(2u)-\sin^2(2u)=\cos(4u)\]

OpenStudy (anonymous):

Is this the correct relationship you are trying to establish?

OpenStudy (anonymous):

Yes

geerky42 (geerky42):

hmm. you know this identity? \(sin^2x = 1 - cos^2x\) and \(cos^2x = \dfrac{1+cos(2x)}{2}\)

OpenStudy (agent0smith):

Erisu Nakayama, I'd start by working on the right hand side - use the angle reduction formula.

geerky42 (geerky42):

what is angle reduction formula?

OpenStudy (agent0smith):

The one you gave, power reduction formula, same thing.

OpenStudy (anonymous):

so I get [-1+cos(2u)+cos^2(2u) ]/2

geerky42 (geerky42):

actually -1+cos^2(2u)+cos^2(2u) = 2cos^2(2u) - 1

OpenStudy (agent0smith):

use\[\Large \cos^2(2x) = \frac{ 1+\cos(4x) }{ 2 }\] also, she's pretttyyy: http://cdn.wallgig.net/2013/12/31/k0oxed3c_wallpaper_2943511.jpg

geerky42 (geerky42):

now for cos^2(2u), apply second identity i gave you.

OpenStudy (anonymous):

oh I got it thanks!

geerky42 (geerky42):

glad we helped

OpenStudy (anonymous):

so i get -.5 + cos(4x)+.5cos^2(4x)

geerky42 (geerky42):

huh? \(2cos^2(2u) - 1 \\ = 2\left(\dfrac{cos(4u)+1}{2}\right)-1\)

geerky42 (geerky42):

2 cancel each other and 1 and -1 cancel each other so that left you cos(4u)